OpenStudy (speedsterpro):

What value could be added to 2/15 to make the sum greater than 1/2 A. 1/15 B. 8/30 C. 5/15 D. 8/15

OpenStudy (madisongoode):

Just add 2/15 to A,B,C, and D, and you'll get your answer when you get an answer above one half.

OpenStudy (madisongoode):

If you need more help, on adding them. let me know.

OpenStudy (madisongoode):

Okay, so, 1/15 + 2/15 = 1/5 8/30 + 2/15 = 2/5 5/15 + 2/15 = 7/15 8/15 + 2/15 = 2/3

OpenStudy (madisongoode):

So which one is greater than 1/2 ?